# itemsep=1pt,topsep=2pt,parsep=1pt,partopsep=1pt

### v2

@@ 2.6. Coarse graining procedure
(2.37) ${\mathcal{T}}_{N,\varepsilon}(\mathrm{i}):=((\mathrm{i}-1)\varepsilon N,\mathrm{i}\varepsilon N],\quad\mathrm{i}=1,\cdots,\left\lfloor\frac{1}{\varepsilon}\right\rfloor.$
Fix a threshold
(2.38) $K_{\varepsilon}:=\Big{(}\log\frac{1}{\varepsilon}\Big{)}^{6}$
@@
(2.39) $\begin{split}{\mathcal{A}}_{\varepsilon}^{\text{\rm(no triple)}}:=\bigcup\limits_{k\in{\mathbbm{N}}}\Big{\{}&(\mathrm{i}_{1},\cdots,\mathrm{i}_{k})\in{\mathbbm{N}}^{k}:K_{\varepsilon}\leq\mathrm{i}_{1}<\cdots<\mathrm{i}_{k}\leq\left\lfloor\frac{1}{\varepsilon}\right\rfloor-K_{\varepsilon},\\ &\text{such that if}~{}\mathrm{i}_{j+1}-\mathrm{i}_{j}<K_{\varepsilon},\text{then}~{}\mathrm{i}_{j+2}-\mathrm{i}_{j+1}\geq K_{\varepsilon}\Big{\}}.\end{split}$
@@
(2.40) $X_{d,f}:=\begin{cases}\zeta_{d},&\quad\text{if}~{}d=f,\\ \zeta_{d}\widetilde{Z}_{d,f}^{\omega,\beta_{N}}(0,0)\zeta_{f},&\quad\text{if}~{}f\geq d+1,\end{cases}$
@@
(2.41) $\begin{split}{\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(no triple)}}(\varphi,\psi):=q_{0,N}^{N}(\varphi,\psi)+&\frac{1}{\sqrt{N}}\sum\limits_{k=1}^{\left(\log\frac{1}{\varepsilon}\right)^{2}}\sum\limits_{(\mathrm{i}_{1},\cdots,\mathrm{i}_{k})\in{\mathcal{A}}_{\varepsilon}^{\text{\rm(no triple)}}}\sum\limits_{d_{1}\leq f_{1}\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i}_{1}),\cdots,d_{k}\leq f_{k}\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i}_{k})}\\ &q_{0,d_{1}}^{N}(\varphi,0)X_{d_{1},f_{1}}\bigg{\{}\prod\limits_{j=2}^{k}p_{d_{j}-f_{j-1}}(0)X_{d_{j},f_{j}}\bigg{\}}q_{f_{k},N}^{N}(0,\psi),\end{split}$
where we rule out the consecutive visits to mesoscopic intervals ${\mathcal{T}}_{N,\varepsilon}(\mathrm{i}_{j}),{\mathcal{T}}_{N,\varepsilon}(\mathrm{i}_{j+1})$ and ${\mathcal{T}}_{N,\varepsilon}(\mathrm{i}_{j+2})$ with $\mathrm{i}_{j+1}-\mathrm{i}_{j}<K_{\varepsilon}$ and $\mathrm{i}_{j+2}-\mathrm{i}_{j+1}<K_{\varepsilon}$.
We will approximate $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$ by ${\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(no triple)}}(\varphi,\psi)$ in $L_{2}$.
@@
(2.42) $\begin{split}\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}:=\bigcup\limits_{r\in{\mathbbm{N}}}\Big{\{}&(\vec{\mathrm{i}}_{1},\cdots,\vec{\mathrm{i}}_{r})\in{\mathbbm{N}}^{r}:K_{\varepsilon}\leq\vec{\mathrm{i}}_{1}<\cdots<\vec{\mathrm{i}}_{k}\leq\Big{\lfloor}\frac{1}{\varepsilon}\Big{\rfloor}-K_{\varepsilon},~{}\text{such that}\\ &|\vec{\mathrm{i}}_{j}|<K_{\varepsilon},\forall~{}j=1,\cdots,r,~{}{\rm dist}(\vec{\mathrm{i}}_{j-1},\vec{\mathrm{i}}_{j})\geq K_{\varepsilon},\forall~{}j=2,\cdots,r\Big{\}}.\end{split}$
@@
(2.43) $\Theta(\vec{\mathrm{i}})=\Theta_{N,\varepsilon}(\vec{\mathrm{i}}):=\begin{cases}\frac{1}{\sqrt{\varepsilon N}}\sum\limits_{d\leq f\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i})}X_{d,f},&\text{if}~{}|\vec{\mathrm{i}}|=1,\\ \frac{1}{\sqrt{\varepsilon N}}\sum\limits_{d\leq f\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i})}\sum\limits_{d^{\prime}\leq f^{\prime}\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i}^{\prime})}\big{(}X_{d,f}\times p_{d^{\prime}-f}(0)\times X_{d^{\prime},f^{\prime}}\big{)},\quad&\text{if}~{}|\vec{\mathrm{i}}|>1.\end{cases}$
@@
(2.44) $\begin{split}{\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi):=g_{1}(\varphi,\psi)+&\sqrt{\varepsilon}\sum\limits_{r=1}^{(\log\frac{1}{\varepsilon})^{2}}\sum\limits_{(\vec{\mathrm{i}}_{1},\cdots,\vec{\mathrm{i}}_{r})\in\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}}\sum\limits_{\mathrm{a},\mathrm{b}\in{\mathbbm{Z}}}\\ &\varphi_{N,\varepsilon}(\mathrm{a})g_{\mathrm{i}_{1}}(\mathrm{a})\Theta(\vec{\mathrm{i}}_{1})\times\bigg{\{}\prod_{j=2}^{r}g_{(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}(0)\Theta(\vec{\mathrm{i}}_{j})\bigg{\}}g_{(\lfloor\frac{1}{\varepsilon}\rfloor-\mathrm{i}^{\prime}_{r})}(\mathrm{b})\psi_{N,\varepsilon}(\mathrm{b}),\end{split}$
@@
(2.45) $\varphi_{N,\varepsilon}(\mathrm{a}):=\frac{1}{\sqrt{\varepsilon N}}\sum\limits_{u\in(\mathrm{a}\sqrt{\varepsilon N},(\mathrm{a}+1)\sqrt{\varepsilon N}]\cap{\mathbbm{Z}}}\varphi_{N}(u),\quad\psi_{N,\varepsilon}(\mathrm{b}):=\frac{1}{\sqrt{\varepsilon N}}\sum\limits_{v\in(\mathrm{b}\sqrt{\varepsilon N},(\mathrm{b}+1)\sqrt{\varepsilon N}]\cap{\mathbbm{Z}}}\varphi_{N}(v).$
@@
(2.46) $\begin{split}\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta):=g_{1}(\varphi,\psi)+&\sqrt{\varepsilon}\sum\limits_{r=1}^{(\log\frac{1}{\varepsilon})^{2}}\sum\limits_{(\vec{\mathrm{i}}_{1},\cdots,\vec{\mathrm{i}}_{r})\in\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}}\sum\limits_{\mathrm{a},\mathrm{b}\in{\mathbbm{Z}}}\\ &\varphi_{\varepsilon}(\mathrm{a})g_{\mathrm{i}_{1}}(\mathrm{a})\Theta(\vec{\mathrm{i}}_{1})\times\bigg{\{}\prod_{j=2}^{r}g_{(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}(0)\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{j})\bigg{\}}g_{(\lfloor\frac{1}{\varepsilon}\rfloor-\mathrm{i}^{\prime}_{r})}(\mathrm{b})\psi_{\varepsilon}(\mathrm{b}),\end{split}$
@@
(2.47) $\varphi_{\varepsilon}(\mathrm{a}):=\int_{\mathrm{a}}^{\mathrm{a}+1}\varphi(\sqrt{\varepsilon}x){\mathrm{d}}x,\quad\psi_{\varepsilon}(\mathrm{b}):=\int_{\mathrm{b}}^{\mathrm{b}+1}\psi(\sqrt{\varepsilon}y){\mathrm{d}}y,\quad\text{for}~{}\mathrm{a},\mathrm{b}\in{\mathbbm{Z}}.$
We will approximate ${\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi)$ by $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta)$ in $L_{2}$. For $u_{0,1}^{\delta}(\varphi,\psi)$, we will also approximate it by $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta_{\delta,\varepsilon})$, where $\Theta_{\delta,\varepsilon}$ is a continuum version of coarse-grained variables (see (8.23)).
2.7. Organizations of the paper
The rest of the paper is organized as follows. In Section 3, we complete the proof of Theorem 1.6. In Section 4, we estimate the moments of $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$. In Section 5, we approximate $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$ by the coarse-grained model. In Section 6, we bound the fourth moment of the coarse-grained model, where a first step is to bound the moments of the coarse-grained disorders. In Section 7, we prove Theorem 1.7 by showing the convergence of the coarse-grained model, where an enhanced Lindeberg principle is applied. In Section 8, we prove Theorem 1.10.
Since Theorem 1.6 proved in Section 3 is novel, and the continuous setting treated in Section 8 was not addressed in [CSZ23], we provide enough details for both sections. In contrast, the proofs in Sections 4-7 are adapted from [CSZ23]. Section 7 is written with full details as the corresponding part in [CSZ23] is only outlined. For the remaining three sections, we explain the strategies in detail but only sketch some computations, so that the paper is convincing while maintaining in a reasonable length. We also point out the major differences and modifications between [CSZ23] and our paper. To make the paper self-contained, we include the enhanced Lindeberg principle in Appendix A.
2.8. Notations
Due to various notations in our paper, we collect some important ones here.
@@ 2.8. Notations
$\bullet$ $X_{d,f}$: The directed polymer partition function $\widetilde{Z}_{d,f}^{\omega,\beta_{N}}(0,0)$ equipped with $\zeta_{d}$ and $\zeta_{f}$, see (2.40).
@@
3. Proof of Theorem 1.6
We are going to show the equivalence between (i) and (ii) in Theorem 1.6. From Subsection 2.1, it suffices to work on $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$ and ${\mathcal{Z}}_{N}^{\beta_{N}}(f,h)$ (see (2.9) and (2.10)), which can be expressed by
@@
$+\sum\limits_{m=1}^{N-1}\sum\limits_{n=m}^{N-1}$ $\Big{(}\int_{{\mathbbm{R}}}\varphi(x)\sqrt{N}q_{x_{N}}(m){\mathrm{d}}x\Big{)}\Big{(}\sqrt{N}Z^{\omega,\beta_{N}}_{m,n}\Big{)}\Big{(}\int_{{\mathbbm{R}}}\psi(x)\sqrt{N}q_{-y_{N}}(N-n){\mathrm{d}}y\Big{)},$
(3.2) ${\mathcal{Z}}_{N}^{\beta_{N}}(f,h)=$ $\sum\limits_{m=0}^{N-1}\sum\limits_{n=m}^{N-1}\Big{(}\int_{\frac{m}{N}}^{\frac{m+1}{N}}f(s){\mathrm{d}}s\Big{)}\Big{(}\sqrt{N}Z_{m,n}^{\omega,\beta_{N}}\Big{)}\Big{(}\int_{\frac{n}{N}}^{\frac{n+1}{N}}h(t){\mathrm{d}}t\Big{)}.$
In particular, it suffices to show that
(a) the first term in (3.1) converges;
(b) the term refers to $m=0$, that is
(3.3) $\mathscr{R}_{N}:=\mathscr{R}_{N}(f,h)=\sqrt{N}\int_{0}^{\frac{1}{N}}f(s){\mathrm{d}}s\sum_{n=0}^{N-1}Z_{0,n}^{\omega,\beta_{N}}\int_{\frac{n}{N}}^{\frac{n+1}{N}}h(t){\mathrm{d}}t,$
converges to 00.
(c) two pairs of integrals,
(3.4) $\int_{{\mathbbm{R}}}\varphi(x)\sqrt{N}q_{x_{N}}(m){\mathrm{d}}x~{}~{}\text{with}~{}~{}\int_{\frac{m}{N}}^{\frac{m+1}{N}}f(s){\mathrm{d}}s\quad\text{and}\quad\int_{{\mathbbm{R}}}\psi(x)\sqrt{N}q_{-y_{N}}(N-n){\mathrm{d}}y~{}~{}\text{with}~{}~{}\int_{\frac{n}{N}}^{\frac{n+1}{N}}h(t){\mathrm{d}}t$
can be made arbitrarily close and uniformly in $N$.
We start by treating Part (a) and Part (b), which are common parts in both (ii)$\Longrightarrow$(i) and (i)$\Longrightarrow$(ii). A key ingredient in the proofs is the local limit theorem [Uch11, Corollary 1.1] for first hitting times. We have that as $N\to\infty$, uniformly in $x\in\frac{1}{\sqrt{N}}{\mathbbm{Z}}$ and $s\in\frac{1}{N}{\mathbbm{N}}$,
(3.5) $q(x\sqrt{N},sN)=\frac{|x|}{\sqrt{2\pi}s^{3/2}N}e^{-\frac{x^{2}}{2s}}+o\Big{(}\frac{|x|\sqrt{N}\vee 1}{(sN)^{3/2}}\wedge\frac{1}{x^{2}N\vee 1}\Big{)}=:\frac{1}{N}Q(x,s)+\frac{1}{N}R_{N}(x,s).$
@@
Hence, to prove Part (a), it suffices to prove that the expectation of the second line of (3.1), that is,
(3.6) $\sum\limits_{m=1}^{N-1}\sum\limits_{n=m}^{N-1}\Big{(}\int_{{\mathbbm{R}}}\varphi(x)\sqrt{N}q_{x_{N}}(m){\mathrm{d}}x\Big{)}\Big{(}\sqrt{N}p_{m,n}(0)\Big{)}\Big{(}\int_{{\mathbbm{R}}}\psi(x)\sqrt{N}q_{-y_{N}}(N-n){\mathrm{d}}y\Big{)},$

### v3

@@ 2.6 Coarse graining procedure
${\mathcal{T}}_{N,\varepsilon}(\mathrm{i}):=((\mathrm{i}-1)\varepsilon N,\mathrm{i}\varepsilon N],\quad\mathrm{i}=1,\cdots,\left\lfloor\frac{1}{\varepsilon}\right\rfloor.$ (2.37)
For $\kappa>0$ in (1.10), fix a threshold
$K_{\varepsilon}:=\Big(\log\frac{1}{\varepsilon}\Big)^{\frac{10}{\kappa\wedge 1}}$ (2.38)
@@
$\begin{split}{\mathcal{A}}_{\varepsilon}^{\text{\rm(no triple)}}:=\bigcup\limits_{k\in{\mathbbm{N}}}\Big\{&(\mathrm{i}_{1},\cdots,\mathrm{i}_{k})\in{\mathbbm{N}}^{k}:K_{\varepsilon}\leq\mathrm{i}_{1}<\cdots<\mathrm{i}_{k}\leq\left\lfloor\frac{1}{\varepsilon}\right\rfloor-K_{\varepsilon},\\ &\text{such that if}~\mathrm{i}_{j+1}-\mathrm{i}_{j}<K_{\varepsilon},\text{then}~\mathrm{i}_{j+2}-\mathrm{i}_{j+1}\geq K_{\varepsilon}\Big\}.\end{split}$ (2.39)
@@
$X_{d,f}:=\begin{cases}\zeta_{d},&\quad\text{if}~d=f,\\ \zeta_{d}\widetilde{Z}_{d,f}^{\omega,\beta_{N}}(0,0)\zeta_{f},&\quad\text{if}~f\geq d+1,\end{cases}$ (2.40)
@@
$\begin{split}{\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(no triple)}}(\varphi,\psi):=q_{0,N}^{N}(\varphi,\psi)+&\frac{1}{\sqrt{N}}\sum\limits_{k=1}^{\left(\log\frac{1}{\varepsilon}\right)^{2}}\sum\limits_{(\mathrm{i}_{1},\cdots,\mathrm{i}_{k})\in{\mathcal{A}}_{\varepsilon}^{\text{\rm(no triple)}}}\sum\limits_{d_{1}\leq f_{1}\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i}_{1}),\cdots,d_{k}\leq f_{k}\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i}_{k})}\\ &q_{0,d_{1}}^{N}(\varphi,0)X_{d_{1},f_{1}}\bigg\{\prod\limits_{j=2}^{k}p_{d_{j}-f_{j-1}}(0)X_{d_{j},f_{j}}\bigg\}q_{f_{k},N}^{N}(0,\psi),\end{split}$ (2.41)
where we rule out the consecutive visits to mesoscopic intervals ${\mathcal{T}}_{N,\varepsilon}(\mathrm{i}_{j}),{\mathcal{T}}_{N,\varepsilon}(\mathrm{i}_{j+1})$ and ${\mathcal{T}}_{N,\varepsilon}(\mathrm{i}_{j+2})$ with $\mathrm{i}_{j+1}-\mathrm{i}_{j}<K_{\varepsilon}$ and $\mathrm{i}_{j+2}-\mathrm{i}_{j+1}<K_{\varepsilon}$. We will approximate $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$ by ${\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(no triple)}}(\varphi,\psi)$ in $L_{2}$.
@@
$\begin{split}\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}:=\bigcup\limits_{r\in{\mathbbm{N}}}\Big\{&(\vec{\mathrm{i}}_{1},\cdots,\vec{\mathrm{i}}_{r})\in{\mathbbm{N}}^{r}:K_{\varepsilon}\leq\vec{\mathrm{i}}_{1}<\cdots<\vec{\mathrm{i}}_{k}\leq\Big\lfloor\frac{1}{\varepsilon}\Big\rfloor-K_{\varepsilon},~\text{such that}\\ &|\vec{\mathrm{i}}_{j}|<K_{\varepsilon},\forall~j=1,\cdots,r,~{\rm dist}(\vec{\mathrm{i}}_{j-1},\vec{\mathrm{i}}_{j})\geq K_{\varepsilon},\forall~j=2,\cdots,r\Big\}.\end{split}$ (2.42)
@@
$\Theta(\vec{\mathrm{i}})=\Theta_{N,\varepsilon}(\vec{\mathrm{i}}):=\begin{cases}\frac{1}{\sqrt{\varepsilon N}}\sum\limits_{d\leq f\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i})}X_{d,f},&\text{if}~|\vec{\mathrm{i}}|=1,\\ \frac{1}{\sqrt{\varepsilon N}}\sum\limits_{d\leq f\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i})}\sum\limits_{d^{\prime}\leq f^{\prime}\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i}^{\prime})}\big(X_{d,f}\times p_{d^{\prime}-f}(0)\times X_{d^{\prime},f^{\prime}}\big),~~&\text{if}~|\vec{\mathrm{i}}|>1.\end{cases}$ (2.43)
@@
$\begin{split}{\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(cg)}}(\varphi,&\psi):=g_{1}(\varphi,\psi)+\sqrt{\varepsilon}\sum\limits_{r=1}^{(\log\frac{1}{\varepsilon})^{2}}\sum\limits_{(\vec{\mathrm{i}}_{1},\cdots,\vec{\mathrm{i}}_{r})\in\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}}\sum\limits_{\mathrm{a},\mathrm{b}\in{\mathbbm{Z}}}\\ &\varphi_{N,\varepsilon}(\mathrm{a})g_{\mathrm{i}_{1}}(\mathrm{a})\Theta(\vec{\mathrm{i}}_{1})\times\bigg\{\prod_{j=2}^{r}g_{(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}(0)\Theta(\vec{\mathrm{i}}_{j})\bigg\}g_{(\lfloor\frac{1}{\varepsilon}\rfloor-\mathrm{i}^{\prime}_{r})}(\mathrm{b})\psi_{N,\varepsilon}(\mathrm{b}),\end{split}$ (2.44)
@@
$\begin{split}\varphi_{N,\varepsilon}(\mathrm{a})&:=\frac{1}{\sqrt{\varepsilon N}}\sum\limits_{u\in(\mathrm{a}\sqrt{\varepsilon N},(\mathrm{a}+1)\sqrt{\varepsilon N}]\cap{\mathbbm{Z}}}\varphi_{N}(u),\\ \quad\psi_{N,\varepsilon}(\mathrm{b})&:=\frac{1}{\sqrt{\varepsilon N}}\sum\limits_{v\in(\mathrm{b}\sqrt{\varepsilon N},(\mathrm{b}+1)\sqrt{\varepsilon N}]\cap{\mathbbm{Z}}}\varphi_{N}(v).\end{split}$ (2.45)
@@
$\begin{split}\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|&\Theta)=\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta_{N,\varepsilon}):=g_{1}(\varphi,\psi)+\sqrt{\varepsilon}\sum\limits_{r=1}^{(\log\frac{1}{\varepsilon})^{2}}\sum\limits_{(\vec{\mathrm{i}}_{1},\cdots,\vec{\mathrm{i}}_{r})\in\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}}\sum\limits_{\mathrm{a},\mathrm{b}\in{\mathbbm{Z}}}\\ &\varphi_{\varepsilon}(\mathrm{a})g_{\mathrm{i}_{1}}(\mathrm{a})\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{1})\times\bigg\{\prod_{j=2}^{r}g_{(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}(0)\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{j})\bigg\}g_{(\lfloor\frac{1}{\varepsilon}\rfloor-\mathrm{i}^{\prime}_{r})}(\mathrm{b})\psi_{\varepsilon}(\mathrm{b}),\end{split}$ (2.46)
@@
$\varphi_{\varepsilon}(\mathrm{a}):=\int_{\mathrm{a}}^{\mathrm{a}+1}\varphi(\sqrt{\varepsilon}x){\mathrm{d}}x,\quad\psi_{\varepsilon}(\mathrm{b}):=\int_{\mathrm{b}}^{\mathrm{b}+1}\psi(\sqrt{\varepsilon}y){\mathrm{d}}y,\quad\text{for}~\mathrm{a},\mathrm{b}\in{\mathbbm{Z}}.$ (2.47)
We will approximate ${\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi)$ by $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta)$ in $L_{2}$. For $\iint_{{\mathbbm{R}}^{2}}\!\varphi(x)\widetilde{u}_{0,1}^{\delta}\!({\mathrm{d}}x,{\mathrm{d}}y)\psi(y)$, we will also approximate it by $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta_{\delta,\varepsilon})$, where $\Theta_{\delta,\varepsilon}$ is a continuum version of coarse-grained variables (see (8.23)).
2.7 Organizations of the paper
The rest of the paper is organized as follows. In Section 3, we complete the proof of Theorem 1.6. In Section 4, we estimate the higher moments of $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$. In Section 5, we approximate $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$ by the coarse-grained model. In Section 6, we bound the fourth moment of the coarse-grained model, where a first step is to bound the moments of the coarse-grained disorders. In Section 7, we prove Theorem 1.7 by showing the convergence of the coarse-grained model, where an enhanced Lindeberg principle is applied. In Section 8, we prove Theorem 1.10.
Since Theorem 1.6 proved in Section 3 is novel, and the continuous setting treated in Section 8 was not addressed in [15], we provide enough details for both sections. In contrast, the proofs in Sections 4-7 are adapted from [15]. Section 7 is written with full details as the corresponding part in [15] is only outlined. For the remaining three sections, we explain the strategies in detail but only sketch some computations, so that the paper is convincing while maintaining in a reasonable length. We also point out the major differences and modifications between [15] and our paper. To make the paper self-contained, we include the enhanced Lindeberg principle in Appendix A and some local limit estimates in Appendix B.
2.8 Notations
Here we collect some important notations in our paper:
@@ 2.8 Notations
$\bullet$ $X_{d,f}$: The directed polymer partition function $\widetilde{Z}_{d,f}^{\omega,\beta_{N}}(0,0)$ equipped with boundary disorders $\zeta_{d}$ and $\zeta_{f}$, see (2.40).
@@
3 Proof of Theorem 1.6
We are going to show the equivalence between the two convergences (i) and (ii) in Theorem 1.6. From Subsection 2.1, it suffices to work on $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$ and ${\mathcal{Z}}_{N}^{\beta_{N}}(f,h)$ (see (2.9) and (2.10)), which can be expressed by
$\begin{array}[]{l}\displaystyle\qquad\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)=\iint_{{\mathbbm{R}}\times{\mathbbm{R}}}\varphi(x)\sqrt{N}\hskip 1.0pt{\mathbf{P}}^{x_{N}}(S_{N}=y_{N},\tau_{1}\geq N)\psi(y){\mathrm{d}}x{\mathrm{d}}y\\[5.0pt] \displaystyle\quad+\sum\limits_{m=1}^{N-1}\sum\limits_{n=m}^{N-1}\Big(\int_{{\mathbbm{R}}}\varphi(x)\sqrt{N}q_{x_{N}}(m){\mathrm{d}}x\Big)\Big(\sqrt{N}Z^{\omega,\beta_{N}}_{m,n}\Big)\Big(\int_{{\mathbbm{R}}}\psi(x)\sqrt{N}q_{-y_{N}}(N-n){\mathrm{d}}y\Big),\end{array}$ (3.1)
${\mathcal{Z}}_{N}^{\beta_{N}}(f,h)=\sum\limits_{m=0}^{N-1}\sum\limits_{n=m}^{N-1}\Big(\int_{\frac{m}{N}}^{\frac{m+1}{N}}f(s){\mathrm{d}}s\Big)\Big(\sqrt{N}Z_{m,n}^{\omega,\beta_{N}}\Big)\Big(\int_{\frac{n}{N}}^{\frac{n+1}{N}}h(t){\mathrm{d}}t\Big).$ (3.2)
We need to show that the convergence of (3.2) for all $f,h\in C([0,1])$) implies the convergence of (3.1) for all $\varphi,\psi\in C_{c}({\mathbbm{R}})$, and vice versa. To this end, we need the following ingredients.
Part (a): the first (deterministic) term in the right-hand side of (3.1) converges as $N\to\infty$;
Part (b): the term
$\mathscr{R}_{N}:=\mathscr{R}_{N}(f,h)=\sqrt{N}\int_{0}^{\frac{1}{N}}f(s){\mathrm{d}}s\sum_{n=0}^{N-1}Z_{0,n}^{\omega,\beta_{N}}\int_{\frac{n}{N}}^{\frac{n+1}{N}}h(t){\mathrm{d}}t,$ (3.3)
which refers to the sum $m\leq n\leq N-1$ and $m=0$ in (3.2), converges to 0 in $L_{2}$ as $N\to\infty$;
Part (c): Given $f,h\in C([0,1])$ (Given $\varphi,\psi\in C_{c}({\mathbbm{R}})$ resp.), for any $\varepsilon>0$, we can find $\varphi,\psi\in C_{c}({\mathbbm{R}})$ (we can find $f,h\in C([0,1])$ resp.), such that
$\limsup\limits_{N\to\infty}\max\limits_{1\leq m\leq N}$ $N\bigg|\int_{{\mathbbm{R}}}\varphi(x)\sqrt{N}q_{x_{N}}(m){\mathrm{d}}x-\int_{\frac{m}{N}}^{\frac{m+1}{N}}f(s){\mathrm{d}}s\bigg|\leq\varepsilon,$ (3.4)
$\limsup\limits_{N\to\infty}\max\limits_{1\leq m\leq N}$ $N\bigg|\int_{{\mathbbm{R}}}\psi(x)\sqrt{N}q_{-y_{N}}(N-n){\mathrm{d}}y-\int_{\frac{n}{N}}^{\frac{n+1}{N}}h(t){\mathrm{d}}t\bigg|\leq\varepsilon.$ (3.5)
@@
We need the following refined estimate [37, Corollary 1.1] for the first hitting times of random walks with a finite second moment: as $N\to\infty$, uniformly in $x\in\frac{1}{\sqrt{N}}{\mathbbm{Z}}$ and $s\in\frac{1}{N}{\mathbbm{N}}$,
$q_{x_{N}}(sN)=\frac{|x|}{\sqrt{2\pi}s^{3/2}N}e^{-\frac{x^{2}}{2s}}+o\Big(\frac{|x|\sqrt{N}\!\vee\!1}{(sN)^{3/2}}\!\wedge\!\frac{1}{x^{2}N\!\vee\!1}\Big)=:\!\frac{1}{N}Q(x,s)\!+\!\frac{1}{N}R_{N}(x,s).$ (3.6)
@@
Hence, to prove Part (a), it suffices to prove that the expectation of the double summation in (3.1), that is,
$\sum\limits_{m=1}^{N-1}\sum\limits_{n=m}^{N-1}\Big(\int_{{\mathbbm{R}}}\varphi(x)\sqrt{N}q_{x_{N}}(m){\mathrm{d}}x\Big)\Big(\sqrt{N}p_{m,n}(0)\Big)\Big(\int_{{\mathbbm{R}}}\psi(x)\sqrt{N}q_{-y_{N}}(N-n){\mathrm{d}}y\Big),$ (3.7)